OpenStudy (anonymous):

Why the derivative of y= 3/x is equal to -3/x^2. Can some one explain me ?

jimthompson5910 (jim_thompson5910):

3/x is the same as 3x^(-1)

jimthompson5910 (jim_thompson5910):

from there, you use the rule that d/dx[ x^n ] = n*x^(n-1) so that explains why d/dx[ 3x^(-1) ] = 3*d/dx[ x^(-1) ] d/dx[ 3x^(-1) ] = 3*(-1)*x^(-2) d/dx[ 3x^(-1) ] = -3*x^(-2) d/dx[ 3x^(-1) ] = -3/(x^2)

OpenStudy (anonymous):

Thankkkksssss !!1

jimthompson5910 (jim_thompson5910):

you're welcome